# Steering LLMs for Formal Theorem Proving

### v4

@@ 5 Conclusion
Our work presents a deep analysis of the underlying mechanisms that drive natural language thoughts in Large Language Models for formal theorem proving. By isolating and manipulating specific activation directions associated with natural language ”thoughts”, we have demonstrated a method to enhance LLMs’ capabilities in mathematical theorem proving without requiring costly fine-tuning. Our experiments showcase how language models comprehend natural language very differently compared to formal proof steps and theorems. We explore the use of activation vectors to represent the task of producing NL information (or thought) with the proof step. And it performs consistently better than base models trained for theorem proving. We provide a closer look into how language models represent theorems in activation space to generate proofs. The use of steering vectors to isolate layers responsible for formal reasoning shows the promise in this approach. We believe this is a strong step towards understanding why language models can not reason on olympiad level problems natively.
@@ A.2.1 How does Steering help in the proof?
Example: mathd_algebra_156 in miniF2F benchmark.
@@
theorem mathd_algebra_156 (x y : $\mathbb{R}$) (f g : $\mathbb{R}$rightarrow $\mathbb{R}$) (h0 : $\forall$ t, f t = t ^ 4)
@@
$+$ b/mathd_algebra_156_internlm2-5_wo_steering.lean
@@
@@ -11,56 +9,47 @@ theorem mathd_algebra_156 (x y : $\mathbb{R}$) (f g : $\mathbb{R}$rightarrow $\mathbb{R}$) (h0 : $\forall$ t, f t = t
@@ A.2.2 Does Steering introduce only useful natural language information?
Example of mathd_algebra_313 where the intermediate step is incorrectly annotated.
@@
Lean Proof for mathd_algebra_313
@@ A.3 Transferability of Steering Vectors
Problem Statement for Putnam_1988_B1

### v5

@@ 6 Conclusion
Our work presents a deep analysis of the underlying mechanisms that drive natural language thoughts in Large Language Models for formal theorem proving. By isolating and manipulating specific activation directions associated with natural language “thoughts”, we have demonstrated a method to enhance LLMs’ capabilities in mathematical theorem proving without requiring costly fine-tuning. Our experiments showcase how language models comprehend natural language very differently compared to formal proof steps and theorems. We explore the use of activation vectors to represent the task of producing NL information (or thought) with the proof step. And it performs consistently better than base models trained for theorem proving. We provide a closer look into how language models represent theorems in activation space to generate proofs. The use of steering vectors to isolate layers responsible for formal reasoning shows the promise in this approach. We believe this provides insight into the challenges LLMs face on Olympiad-level problems and how activation steering may mitigate them.
@@ A.2.1 How does Steering help in the proof?
Example: Mathd Algebra 156 in miniF2F benchmark.
@@
theorem mathd_algebra_156 (x y : $\mathbb{R}$) (f g : $\mathbb{R}$ $\rightarrow$ $\mathbb{R}$) (h0 : $\forall$ t, f t = t ^ 4)
@@
+++ b/mathd_algebra_156_internlm2-5_wo_steering.lean
@@
@@ -11,56 +9,47 @@ theorem mathd_algebra_156 (x y : $\mathbb{R}$) (f g : $\mathbb{R}$ $\rightarrow$ $\mathbb{R}$) (h0 : $\forall$ t, f t = t
@@ A.2.2 Does Steering introduce only useful natural language information?
Example of Mathd Algebra 313 where the intermediate step is incorrectly annotated.
@@
Lean Proof for Mathd Algebra 313
@@ A.3 Transferability of Steering Vectors
Problem Statement for Putnam 1988 B1